The equation V=12000(0.8)^t describes the value (V) of a car t years after purchase .
a) What was the cars initial value?
b) What was the rate of depreciation expressed as a percentage?
c) What will be the car's value 5 years after it was purchased?
d) Use the graphing calculator to determine the number of years it would take for the car to be worth $824.63 ? Can this part be done without a graphing calculator because I don't have one ?
For (a) I got $12000 , (b) I got 20% (c) I got #3832.16 Please, can someone help me ?

Answer by stanbon(75887)

The equation V=12000(0.8)^t describes the value (V) of a car t years after purchase .
a) What was the cars initial value?
Find t(0) to get the initial value:
V(0) = 12000*1 = $12,000
--------------------------
b) What was the rate of depreciation expressed as a percentage?
Each year you are taking 0.80 of the previous years value.
The depreciation is 0.20 or 20%
------------------
c) What will be the car's value 5 years after it was purchased?
Find V(5)
---------------
d) Use the graphing calculator to determine the number of years it would take for the car to be worth $824.63 ? Can this part be done without a graphing
calculator because I don't have one ?
-------
Let V(t) = 12000(0.8)^t and solve for "t"
I got t = 12
---------------------
For (a) I got $12000
Good
(b) I got 20%
Good
(c) I got #3832.16
I got $3,932.20
===============
Cheers,
Stan H.